The power of the uv radiation emitted by the sun is 3.79e25 w. What is the power of this uv radiation that is incident on the eArth.? The earth is 1.5e11 m from the sun and the cross-sectional area of the earth exposed to the suns light is 1.3e14m^2. Assume that the suns radiation is emitted uniformly in all directions. The area of a sphere is 4 pir^2
1
Expert's answer
2020-04-29T09:49:47-0400
"P_E=P\frac{A_E}{4\pi r^2}"

"P_E=(3.79\cdot 10^{25})\frac{(1.3\cdot 10^{14})}{4\pi (1.5\cdot 10^{11})^2}=(1.7\cdot 10^{16})\ W"